# Derived deformation functors, Koszul duality, and Maurer–Cartan spaces

### v1

@@ 1. Derived deformation functors
Where classical deformation theory concerns functors on local Artinian rings, derived deformation theory looks at functors on enhancements such as differential graded or simplicial rings. We will focus on the former, which only apply in characteristic 00. Fix a field $k$ of characteristic 00.
@@ 1.1. Artinian cdgas and DGLAs
Write $dg\mathrm{Art}_{k}$ for the category of local Artinian $k$-cdgas, $dg_{+}\mathrm{Art}_{k}\subset dg\mathrm{Art}$ for the full subcategory of non-negatively graded objects $\ldots A_{2}\to A_{1}\to A_{0}$, and $\mathrm{Art}_{k}\subset dg_{+}\mathrm{Art}_{k}$ for the full subcategory on objects concentrated in degree 00.
@@
As is traditional in derived deformation theory [Hin98, Man99b, Pri07], the nilpotence condition in Definition 1.1 is strict: there exists $n>0$ such that any product of $n$ elements in $\mathfrak{m}(A_{\bullet})$ is 00. The later formulation of [Lur11] instead considers cdgas which are only homologically Artinian, but they break the Maurer–Cartan functor which will be our main focus, although their homotopy theory is equivalent by [Pri10, Proposition 2.7] or [Boo20, Corollary 4.4.4].
@@ 1.2. Extended functors associated to DGLAs (after [Man99b, Kon94] …)
Denote the quotient set $\mathrm{MC}(L)/\mathrm{Gg}(L)$ by $\mathrm{Def}(L)$, and the quotient groupoid $[\mathrm{MC}(L)/\mathrm{Gg}(L)]$ by $\mathfrak{Del}(L)$ (the Deligne groupoid).
@@
of groups and sets (where $L$ is assumed nilpotent for the last two sequences).
@@
Given $\omega\in\mathrm{MC}(M)$, surjectivity of $L^{1}\to M^{1}$ allows us to choose a lift $\tilde{\omega}\in L^{1}$, and then $\kappa(\tilde{\omega}):=d\tilde{\omega}+\frac{1}{2}[\tilde{\omega},\tilde{\omega}]$ lies in $\ker(L^{2}\to M^{2})=I^{2}$. Moreover, since $[\tilde{\omega},[\tilde{\omega},\tilde{\omega}]]=0$ and $d^{2}=0$, we have $d\kappa(\tilde{\omega})=[d\tilde{\omega},\tilde{\omega}]=[\kappa(\tilde{\omega}),\tilde{\omega}],$ which vanishes because $I$ is central. Hence $\kappa(\tilde{\omega})\in\mathrm{Z}^{2}I$. Any other lift for $\omega$ takes the form $\tilde{\omega}+x$ for $x\in I^{1}$, with $\kappa(\tilde{\omega}+x)=\kappa(\tilde{\omega})+dx$ (centrality again). Thus the class $[\kappa(\tilde{\omega})]\in\mathrm{H}^{2}I$ is the potential obstruction to lifting $\omega$ to $\mathrm{MC}(L)$, and the set of lifts is a torsor for $\mathrm{Z}^{1}I$ under addition. The exact sequence for $\mathrm{Gg}$ is immediate, and that for $\mathrm{Def}$ then follows by passing to quotients. ∎
@@
(finite sum because $A$ is Artinian), with bracket $[u\otimes a,v\otimes b]=(-1)^{\deg v\deg a}[u,v]\otimes ab$ and differential $d(u\otimes a)=(du)\otimes a+(-1)^{\deg u}u\otimes da.$
@@
[…]
@@ 1.3. What do these classify?
For $V_{\bullet}$ a chain complex of $k$-vector spaces, we have a DGLA $L:=\underline{\mathrm{End}}(V)$, where $L^{n}=\prod_{i}\mathrm{Hom}_{k}(V_{i+n},V_{i})$, with bracket $[f,g]=f\circ g-(-1)^{\deg f\deg g}g\circ f$ and differential $d(f):=[d,f]$.
Extending $A$-linearly, elements of $(\mathrm{Tot}\,(L\otimes\mathfrak{m}(A)))^{1}$ can be identified with $A$-linear morphisms $\mathrm{Tot}\,(A\otimes V)\to\mathrm{Tot}\,(\mathfrak{m}(A)\otimes V)$ of homological degree $-1$. Such an element satisfies the Maurer–Cartan equation if and only if the corresponding map $\omega\colon\thinspace A\otimes V\to A\otimes V[1]$ satisfies $(d+\omega)\circ(d+\omega)=0$. Meanwhile,
@@
Beware that this flatness condition interacts poorly with quasi-isomorphism, and in particular the tensor product does not necessarily compute the derived tensor product unless both $A$ and $V$ are concentrated in non-negative chain degrees (see [Hin99, Example 4.3] for a counterexample). It is instead a derived tensor product of the second kind as in [Pos09, §3.12].
@@
For algebras $R$ over $k$-linear dg operads $\mathcal{P}$, we can similarly consider the DGLA $L:=\underline{\mathrm{Der}}_{\mathcal{P}}(R,R)$ of derivations of $R$ as a graded $\mathcal{P}$-algebra, and then $\mathrm{MC}(L,A)$ parametrises deformations of the structural derivation $d$ on $R$, i.e. closed derivations $d^{\prime}$ of homological degree $-1$ on the $\mathcal{P}$-algebra $R\otimes_{k}A$ in graded $A$-modules with $d^{\prime}\equiv d\mod\mathfrak{m}(A)$.
@@ 1.4.1. Extended deformation functors
Given a deformation functor $F$, define $\mathrm{H}^{i}(F):=F(k[\epsilon_{n}])$; this is in fact a $k$-vector space.
In this notation, Example 1.8 says that $\mathrm{H}^{i}(\mathrm{Def}(L,-))\cong\mathrm{H}^{i+1}(L)$.
@@ 1.4.2. Obstructions
Let $\tilde{B}$ be the cone of $e\colon\thinspace I\to A$, regarded as a cdga in the obvious way with $I\cdot I=0$. Then we have an acyclic small extension $\phi\colon\thinspace\tilde{B}\to B$ and a natural surjection $\rho\colon\thinspace\tilde{B}\twoheadrightarrow k\oplus I[1]$ such that $A=\tilde{B}\times_{\rho,k\oplus I[1]}k$.
@@ 2.1. Pro-Artinian cdgas, dg coalgebras and the bar/cobar construction
Since $dg\mathrm{Art}_{k}$ is an Artinian category, by [Gro60, Corollary to Proposition 3], we can assume that all morphisms in our inverse systems are surjections and we can contravariantly identify $\mathrm{pro}(dg\mathrm{Art}_{k})$ with the category $\mathrm{lex}(dg\mathrm{Art}_{k},\mathrm{Set})$ of left exact (i.e. finite limit-preserving) set-valued functors on $dg\mathrm{Art}_{k}$.
@@
Beware that $\beta^{*}$-quasi-isomorphism is a more restrictive notion than quasi-isomorphism, although they agree for objects of $\mathrm{pro}(dg_{+}\mathrm{Art}_{k})\subset\mathrm{pro}(dg_{+}\mathrm{Art}_{k})$ by [Hin98, Proposition 3.3.2]. The final statement of the proposition implies that the homotopy category $\mathrm{Ho}(\mathrm{pro}(dg\mathrm{Art}_{k}))$ is given by localising $\mathrm{pro}(dg\mathrm{Art}_{k})$ at transfinite composites of acyclic small extensions. This homotopy category is a non-abelian analogue of the derived categories of the second kind we encountered in Example 1.9.
@@
There is nothing particularly special about Lie and commutative algebras in the formation of this adjunction. There are similar constructions for algebras over any Koszul dual pair of dg operads [^3] This question inspired the notion of Koszul duality for operads, first proposed in [Dri88]., with the Maurer–Cartan functor still providing the adjunction because tensor product of algebras of the respective types yields a DGLA. Such equivalences are explicitly described in a slightly different setting in [CCN22], but also see §55.1 below.
@@ 2.2. Functors on $\mathrm{pro}(dg\mathrm{Art}_{k})$
The left exact functor $\mathrm{Gg}\times\mathrm{MC}$ is a path object in $\mathrm{lex}(dg\mathrm{Art}_{k},\mathrm{Set})$ for the fibrant object $\mathrm{MC}$.
@@
Lemma 1.5 implies that $\mathrm{MC}$ is fibrant, so to show that $\mathrm{Gg}\times\mathrm{MC}$ is a path object, it suffices to show that the map $\omega\mapsto(\mathrm{id},\omega)$ from $\mathrm{MC}$ is a weak equivalence and that the map $\mathrm{Gg}\times\mathrm{MC}\to\mathrm{MC}\times\mathrm{MC}$ sending $(g,\omega)$ to $(g\star\omega,\omega)$ is a fibration.
For the first property, just observe that for any small extension $A\to B$, the map $\mathrm{Gg}(A)\to\mathrm{Gg}(B)$ is surjective, so the projection map $\mathrm{Gg}\times\mathrm{MC}\to\mathrm{MC}$ is a trivial fibration.
For the second property, we need to show that for any small extension $\phi\colon\thinspace A\to B$ with kernel $I$ and any element $(g,\omega,\omega^{\prime})$ in $\mathrm{Gg}(B)\times\mathrm{MC}(A)^{2}$ such that $g\star\phi(\omega)=\phi(\omega^{\prime})$, there exists an element $\tilde{g}\in\mathrm{Gg}(A)$ lifting $g$ with $\tilde{g}\star\omega=\omega^{\prime}$. To see this, lift $g$ to an element $\breve{g}\in\mathrm{Gg}(A)$ and note that $\breve{g}\star\omega-\omega^{\prime}\in\mathrm{Z}^{1}I$. Since $I$ is acyclic, this equals $dx$ for some $x\in I^{0}$, and then setting $\tilde{g}:=\breve{g}-x$ gives the required element. ∎
@@ 3. Simplicial functors
Furthermore, functors such as those in Examples 1.9 and 1.10 only govern derived deformations when restricted to $dg_{+}\mathrm{Art}\subset dg\mathrm{Art}$, but restricting the functor loses information such as negative cohomology groups of the DGLA. By working with simplicial set-valued functors, we can safely restrict to $dg_{+}\mathrm{Art}$ without losing information.

### v2

# Derived deformation functors, Koszul duality, and Maurer-Cartan spaces

@@ 1. Derived deformation functors
Where classical deformation theory concerns functors on local Artinian rings, derived deformation theory looks at functors on enhancements such as differential graded or simplicial rings. We will focus on the former, which only apply in characteristic 0. Fix a field $k$ of characteristic 0.
@@ 1.1. Artinian cdgas and DGLAs
Write $dg\mathrm{Art}_{k}$ for the category of local Artinian $k$-cdgas, $dg_{+}\mathrm{Art}_{k}\subset dg\mathrm{Art}$ for the full subcategory of non-negatively graded objects $\ldots A_{2}\to A_{1}\to A_{0}$, and $\mathrm{Art}_{k}\subset dg_{+}\mathrm{Art}_{k}$ for the full subcategory on objects concentrated in degree 0.
@@
As is traditional in derived deformation theory [Hin98, Man99b, Pri07], the nilpotence condition in Definition 1.1 is strict: there exists $n>0$ such that any product of $n$ elements in $\mathfrak{m}(A_{\bullet})$ is 0. The formulation of [Lur04, §6.2] and [Lur11] instead considers cdgas which are only homologically Artinian, but they break the Maurer–Cartan functor which will be our main focus, although their homotopy theory is equivalent by [Pri10, Proposition 2.7] or [Boo20, Corollary 4.4.4]. [^2] Specifically, the Postnikov tower ensures that homologically Artinian cdgas lie in the closure of $dg_{+}\mathrm{Art}_{k}$ under homotopy pullbacks, so [Boo20, Corollary 4.4.4] implies that derived completion gives a fully faithful $\infty$-functor from them to the simplicial localisation $\mathbf{L}^{\mathcal{W}}\mathrm{pro}(dg_{+}\mathrm{Art}_{k})$ of the pro-category at quasi-isomorphisms $\mathcal{W}$.
If, for $A,B\in dg_{+}\mathrm{Art}_{k}$, we have a span $A\xleftarrow{p}A^{\prime}\xrightarrow{q}B$ in $\mathrm{pro}(dg_{+}\mathrm{Art}_{k})$ with $p$ a trivial fibration and $q$ a fibration, then [Pri07, Lemma 4.5]11footnotemark: 1allows us to write $p$ as a filtered limit of a system of acyclic small extensions in $dg_{+}\mathrm{Art}$. Since $q$ must factor through some term $A^{\prime\prime}$ in the system, we can replace $A^{\prime}$ with its quotient $A^{\prime\prime}\in dg_{+}\mathrm{Art}_{k}$ with the same properties. Thus the essential image of $dg_{+}\mathrm{Art}_{k}\to\mathrm{Ho}(\mathrm{pro}(dg_{+}\mathrm{Art}_{k}))$ is closed under homotopy pullbacks.
Since $\mathcal{W}$-local equivalences (in the sense of left Bousfield localisation) are preserved under filtered colimits, that lemma also implies that the Yoneda functor $h\colon\thinspace\mathrm{pro}(dg_{+}\mathrm{Art}_{k})^{\mathrm{opp}}\hookrightarrow(\mathrm{Set})^{dg_{+}\mathrm{Art}_{k}}\hookrightarrow(s\mathrm{Set})^{dg_{+}\mathrm{Art}_{k}}$ sends quasi-isomorphisms to $\mathcal{W}$-local equivalences. If $\tilde{A}^{\bullet}$ is a cofibrant cosimplicial frame in $\mathrm{pro}(dg_{+}\mathrm{Art}_{k})$ for $A\in dg_{+}\mathrm{Art}_{k}$, the map $h_{A}\to h_{\tilde{A},\bullet}$ is thus a $\mathcal{W}$-local equivalence, and $h_{\tilde{A},\bullet}$ is $\mathcal{W}$-local. Combining [Hov99, §5.4] with the proof of [DK87, Theorem 2.2], it follows that $\mathbf{L}^{\mathcal{W}}dg_{+}\mathrm{Art}_{k}\to\mathbf{L}^{\mathcal{W}}\mathrm{pro}(dg_{+}\mathrm{Art}_{k})$ is also a fully faithful $\infty$-functor.
22footnotetext: Rather, we partially generalise to $\mathrm{pro}(dg_{+}\mathrm{Art}_{k})$. Take a minimal dg ideal $I\leq A^{\prime}$ for which $A^{\prime}/I\to A$ has the desired property; this must exist by pro-Artinianity. Then $\mathrm{H}_{*}I=0$ and any acyclic subcomplex of $I$ containing $\mathfrak{m}(A)I$ would contradict minimality, so $(I/\mathfrak{m}(A)I)_{n}\cong\mathrm{H}_{n}(I/\mathfrak{m}(A)I)\cong\mathrm{H}_{n-1}(\mathfrak{m}(A)I)$. Thus $I_{\leq n-1}=0$ implies $I_{\leq n}=0$, so $I=0$.
@@ 1.2. Extended functors associated to DGLAs (after [Man99b, Kon94] …)
Denote the quotient set $\mathrm{MC}(L)/\mathrm{Gg}(L)$ by $\mathrm{Def}(L)$, and the quotient groupoid $\mathrm{MC}(L)/\mathrm{Gg}(L)]$ by $\mathfrak{Del}(L)$ (the Deligne groupoid).
@@
of groups and sets (where $L^{0}$ is assumed nilpotent for the last two sequences).
@@
Given $\omega\in\mathrm{MC}(M)$, surjectivity of $L^{1}\to M^{1}$ allows us to choose a lift $\tilde{\omega}\in L^{1}$, and then $\kappa(\tilde{\omega}):=d\tilde{\omega}+\frac{1}{2}[\tilde{\omega},\tilde{\omega}$] lies in $\ker(L^{2}\to M^{2})=I^{2}$. Moreover, since $\tilde{\omega},[\tilde{\omega},\tilde{\omega}]]=0$ and $d^{2}=0$, we have $d\kappa(\tilde{\omega})=[d\tilde{\omega},\tilde{\omega}]=[\kappa(\tilde{\omega}),\tilde{\omega}$], which vanishes because $I$ is central. Hence $\kappa(\tilde{\omega})\in\mathrm{Z}^{2}I$. Any other lift for $\omega$ takes the form $\tilde{\omega}+x$ for $x\in I^{1}$, with $\kappa(\tilde{\omega}+x)=\kappa(\tilde{\omega})+dx$ (centrality again). Thus the class $\kappa(\tilde{\omega})]\in\mathrm{H}^{2}I$ is the potential obstruction to lifting $\omega$ to $\mathrm{MC}(L)$, and the set of lifts is a torsor for $\mathrm{Z}^{1}I$ under addition. The exact sequence for $\mathrm{Gg}$ is immediate, and that for $\mathrm{Def}$ then follows by passing to quotients. ∎
@@
(finite sum because $A$ is Artinian), with bracket $u\otimes a,v\otimes b]=(-1)^{\deg v\deg a}[u,v]\otimes ab$ and differential $d(u\otimes a)=(du)\otimes a+(-1)^{\deg u}u\otimes da.$
@@
$\mathrm{Def}(L,k[\epsilon_{n}])\cong\mathrm{coker\,}(L^{n}\epsilon_{n}\xrightarrow{d}\mathrm{Z}^{n+1}(L)\epsilon_{n})=\mathrm{H}^{n+1}(L).$
@@ 1.3. What do these classify?
For $V_{\bullet}$ a chain complex of $k$-vector spaces, we have a DGLA $L:=\underline{\mathrm{End}}(V)$, where $L^{n}=\prod_{i}\mathrm{Hom}_{k}(V_{i+n},V_{i})$, with bracket $f,g]=f\circ g-(-1)^{\deg f\deg g}g\circ f$ and differential $d(f):=[d,f]$.
Extending $A$-linearly, elements of $(\mathrm{Tot}\,(L\otimes\mathfrak{m}(A)))^{1}$ can be identified with $A$-linear morphisms $\mathrm{Tot}\,(A\otimes V)\to\mathrm{Tot}\,(\mathfrak{m}(A)\otimes V)$ of homological degree 1-1. Such an element satisfies the Maurer–Cartan equation if and only if the corresponding map $\omega\colon\thinspace A\otimes V\to A\otimes V[1]$ satisfies $(d+\omega)\circ(d+\omega)=0$. Meanwhile,
@@
Beware that this flatness condition interacts poorly with quasi-isomorphism, and in particular $V^{\prime}\otimes_{A}k$ does not necessarily compute $V^{\prime}\otimes^{\mathbf{L}}_{A}k$ unless both $A$ and $V$ are concentrated in non-negative chain degrees (see [Hin99, Example 4.3] for a counterexample). It is instead a derived tensor product of the second kind as in [Pos09, §3.12]; see [Pri25] for analysis of such deformations.
@@
For algebras $R$ over $k$-linear dg operads $\mathcal{P}$, we can similarly consider the DGLA $L:=\underline{\mathrm{Der}}_{\mathcal{P}}(R,R)$ of derivations of $R$ as a graded $\mathcal{P}$-algebra, and then $\mathrm{MC}(L,A)$ parametrises deformations of the structural derivation $d$ on $R$, i.e. closed derivations $d^{\prime}$ of homological degree 1-1 on the $\mathcal{P}$-algebra $R\otimes_{k}A$ in graded $A$-modules with $d^{\prime}\equiv d\mod\mathfrak{m}(A)$.
@@ 1.4.1. Extended deformation functors
Given a deformation functor $F$, define its $n$th tangent cohomology group by $\mathrm{H}^{n}(F):=F(k[\epsilon_{n}])$; this is in fact a $k$-vector space.
In this notation, Example 1.8 says that $\mathrm{H}^{n}(\mathrm{Def}(L,-))\cong\mathrm{H}^{n+1}(L)$.
@@ 1.4.2. Obstructions
Let $\tilde{B}$ be the cone of $e\colon\thinspace I\to A$, regarded as a cdga in the obvious way with $I\cdot I=0$. Then we have an acyclic small extension $\phi\colon\thinspace\tilde{B}\to B$ and a natural surjection $\rho\colon\thinspace\tilde{B}\twoheadrightarrow k\oplus I[1]$ (with $I[1]$ square-zero) such that $A=\tilde{B}\times_{\rho,k\oplus I[1]}k$.
@@ 2.1. Pro-Artinian cdgas, dg coalgebras and the bar/cobar construction
The category $dg\mathrm{Art}_{k}$ has finite limits and is Artinian in the sense that all objects satisfy the descending chain condition on subobjects. By [Gro60, Corollary to Proposition 3], we can thus assume that all morphisms in our inverse systems are surjections and we can contravariantly identify $\mathrm{pro}(dg\mathrm{Art}_{k})$ with the category $\mathrm{lex}(dg\mathrm{Art}_{k},\mathrm{Set})$ of left exact (i.e. finite limit-preserving) set-valued functors on $dg\mathrm{Art}_{k}$.
@@
Beware that $\beta^{*}$-quasi-isomorphism is a more restrictive notion than quasi-isomorphism, although they agree for objects of $\mathrm{pro}(dg_{+}\mathrm{Art}_{k})\subset\mathrm{pro}(dg\mathrm{Art}_{k})$ by [Hin98, Proposition 3.3.2]. The final statement of the proposition implies that the homotopy category $\mathrm{Ho}(\mathrm{pro}(dg\mathrm{Art}_{k}))$ is given by localising $\mathrm{pro}(dg\mathrm{Art}_{k})$ at limits of filtered systems of acyclic small extensions. This homotopy category is a non-abelian analogue of the derived categories of the second kind we encountered in Example 1.9.
@@
There is nothing particularly special about Lie and commutative algebras in the formation of this adjunction. There are similar constructions for algebras over any Koszul dual pair of dg operads [^5] This question inspired the notion of Koszul duality for operads, first proposed in [Dri88]., with the Maurer–Cartan functor still providing the adjunction because tensor product of algebras of the respective types yields a DGLA. Such equivalences are explicitly described in a slightly different setting in [CCN19], but also see §55.1 below.
@@ 2.2. Functors on $\mathrm{pro}(dg\mathrm{Art}_{k})$
The left exact functor $\mathrm{Gg}(L,-)\times\mathrm{MC}(L,-)$ is a path object in $\mathrm{lex}(dg\mathrm{Art}_{k},\mathrm{Set})$ for the fibrant object $\mathrm{MC}(L,-)$.
@@
Lemma 1.5 implies that $\mathrm{MC}(L,-)$ is fibrant. To show that $\mathrm{Gg}(L,-)\times\mathrm{MC}(L,-)$ is a path object, we need the map $\omega\mapsto(\mathrm{id},\omega)$ from $\mathrm{MC}(L,-)$ to be a weak equivalence and the map $\mathrm{Gg}(L,-)\times\mathrm{MC}(L,-)\to\mathrm{MC}(L,-)\times\mathrm{MC}(L,-)$ sending $(g,\omega)$ to $(g\star\omega,\omega)$ to be a fibration.
For the first property, just observe that for any small extension $A\to B$, the map $\mathrm{Gg}(L,A)\to\mathrm{Gg}(L,B)$ is surjective, so the projection map $\mathrm{Gg}(L,-)\times\mathrm{MC}(L,-)\to\mathrm{MC}(L,-)$ is a trivial fibration.
For the second property, we need to show that for any small extension $\phi\colon\thinspace A\to B$ with kernel $I$ and any element $(g,\omega,\omega^{\prime})$ in $\mathrm{Gg}(L,B)\times\mathrm{MC}(L,A)^{2}$ such that $g\star\phi(\omega)=\phi(\omega^{\prime})$, there exists an element $\tilde{g}\in\mathrm{Gg}(L,A)$ lifting $g$ with $\tilde{g}\star\omega=\omega^{\prime}$. To see this, lift $g$ to an element $\breve{g}\in\mathrm{Gg}(L,A)$ and note that $\breve{g}\star\omega-\omega^{\prime}\in\mathrm{Z}^{1}(L\otimes I)$. Since $I$ is acyclic, this equals $dx$ for some $x\in(L\otimes I)^{0}$, and then setting $\tilde{g}:=\breve{g}-x$ gives the required element. ∎
@@ 3. Simplicial functors
Furthermore, functors such as those in Examples 1.9 and 1.10 only govern derived deformations when restricted to $dg_{+}\mathrm{Art}\subset dg\mathrm{Art}$ (cf. [Pri25]), but restricting the functor loses information such as negative cohomology groups of the DGLA. By working with simplicial set-valued functors, we can safely restrict to $dg_{+}\mathrm{Art}$ without losing information.